In a town of 360 people, 80%can read and write. 90%of the males alone can read and write, but only 180 of the females can read and write. How many females are in the town?
Answer by ptaylor(1894)


Let x=number of females
Then 360-x=number of males
And 0.90(360-x)=number of males that can read and write
0.80(360)= number of males and females that can read and write
Number of males that can read and write + number of females that can read and write equals 0.80(360). So our equation to solve is:
0.9(360-x)+180=0.8(360)
324-0.9x+180=288
504-0.9x=288 subtract 504 from both sides
-0.9x=288-504
-0.9x=-216 divide both sides by -0.9
x=240-------------------------------------number of females
360-x=360-240=120----------------------------number of males
ck
90% of the males = 0.9(120)=108
108+180=288
80% of 360=288

Hope this helps----ptaylor